OpenStudy (anonymous):

write the number in the form a+bi. √-4+10

OpenStudy (anonymous):

Didn't you ask this problem before?

OpenStudy (anonymous):

Or was that someone else asking the same question?

OpenStudy (anonymous):

√-4 + 10 = i√4 + 10 = 10 + 2i